# Recursive PAC-Bayes: A Frequentist Approach to Sequential Prior Updates

### v1

@@ 6 Discussion
We have presented the first PAC-Bayesian bound that supports sequential prior updates and preserves confidence information on the prior. The work closes a long-standing gap between Bayesian and Frequentist learning by making sequential data processing and sequential updates of prior knowledge meaningful and beneficial in the frequentist framework, as it has always been in the Bayesian framework. We have shown that apart from theoretical beauty the approach is beneficial in practice.
@@ Appendix A Illustrations
Figure 1: Evolution of PAC-Bayes. The figure shows how data are used by different methods. Dark yellow shows data used directly for optimization of the indicated quantities. Light yellow shows data involved indirectly through dependence on the prior. Light green shows data used for estimation of the indicated quantities. In Recursive PAC-Bayes data are released and used sequentially chunk-by-chunk, as indicated by the dashed lines. For example, in the $T=4$ case $\mathbb{E}_{\pi_{1}}[L(h)]$ is first evaluated on $S_{1}$ to construct $\pi_{1}$ and $\gamma_{2}$, then in the first recursion step on $S_{1}\cup S_{2}$, in the second step on $S_{1}\cup S_{2}\cup S_{3}$, and in the last step on all $S$.
@@ B.3 Other details in the experiments
The methods in comparisons are trained and evaluated using the procedure described in Section 2 and visually illustrated in Figure 1. We will provide some further details for each method later in the following. For all methods in comparison, we apply the optimization and evaluation method described in Section 5.1. For the approximation described in Section 5.1.1, we set the parameters $c_{1}=c_{2}=3$. The lower bound for the prediction $p_{\min}=1e-5$. The $\delta$ in our bound and all the other methods is selected to be $\delta=0.025$. As mentioned in Section 5.1.2, we use the PAC-Bayes-classic bound by McAllester in replacement of PAC-Bayes-$\operatorname{kl}$ when doing optimization. Note that for all methods, we also have to estimate the empirical loss of the posterior $\mathbb{E}_{\pi}[\cdot]$ described in Section 5.1.3. We also allocate the budget for the union bound for the estimation such that these estimations in the bound are controlled with probability at least $1-\delta^{\prime}$, where we chose $\delta^{\prime}=0.01$. Therefore, the ultimate bounds for all methods hold with probability at least $1-\delta-\delta^{\prime}$. Note that we do not consider such bounds during optimization but only when estimating the bounds.
@@
In the training process of all methods in our experiments, we set the batch size to 250, the number of training epochs to 50, and use stochastic gradient descent with a learning rate of 0.005 and a momentum of 0.95.
@@ B.4.1 Recursive PAC-Bayes
We report the additional results of Recursive PAC-Bayes on MNIST with $T=2$ in Table 4 and $T=4$ in Table 5. We report Recursive PAC-Bayes on Fashion MNIST with $T=2$ in Table 6 and $T=4$ in Table 7.
@@
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .075 (1e-3) .389 (8e-3) .190 (.020)
2 30000 .039 (9e-3) .017 (8e-4) .137 (.012) .331 (.011) .141 (6e-3)
@@
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .079 (8e-4) .399 (.015) .203 (.027)
2 52500 .026 (7e-3) .013 (6e-4) .110 (7e-3) .309 (1e-3) .141 (.011)
3 45000 .052 (3e-3) .004 (2e-4) .096 (3e-3) .251 (3e-3) .107 (5e-3)
4 30000 .050 (1e-3) .003 (3e-4) .089 (2e-3) .215 (4e-3) .104 (5e-3)
@@
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .046 (8e-4) .385 (9e-4) .242 (8e-3)
2 30000 .097 (5e-3) .007 (2e-4) .167 (4e-3) .359 (4e-3) .220 (8e-3)
@@
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t},\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .049 (7e-4) .408 (5e-3) .261 (.017)
2 52500 .071 (4e-3) .007 (3e-4) .141 (3e-3) .345 (2e-3) .190 (4e-3)
3 45000 .089 (4e-3) .002 (2e-4) .127 (4e-3) .299 (5e-3) .187 (6e-3)
4 30000 .091 (7e-3) .002 (1e-4) .129 (7e-3) .279 (6e-3) .193 (7e-3)
@@ B.4.4 Data-informed priors + excess loss
We report the additional results of data-informed priors + excess loss (Mhammedi et al., 2019) on MNIST and Fashion MNIST in Table 10 and 11. As described earlier in Section 2, Section 5, and Section B.3, we evaluate the bound using $S_{2}$ that is independent of the data-informed prior $\pi_{1}$ and the reference prediction rule $h^{*}$. The bound is composed of two parts: a bound on the excess loss of $\rho$ with respect to $h^{*}$ (Excess bound) and a single hypothesis bound on $h^{*}$ ($h^{*}$ bound). We report the two components of the bound in Table 10. We provide further details to compute these bounds from the losses of their corresponding quantities in Table 11.

### v3

@@ 6 Discussion
We have presented the first PAC-Bayesian bound that supports sequential prior updates and preserves confidence information on the prior. The work closes a long-standing gap between Bayesian and Frequentist learning by making sequential data processing and sequential updates of prior knowledge meaningful and beneficial in the frequentist framework, as it has always been in the Bayesian framework. We have shown that apart from theoretical beauty the approach is highly beneficial in practice.
The Recursive PAC-Bayes framework is extremely rich and powerful, and leads to numerous directions for future research, some of which we briefly sketch next.
  • The decomposition in (2) applies to any loss function, including unbounded losses. It would be interesting to find additional applications to it.
  • While we have restricted ourselves to the zero-one loss function to illustrate the use of PAC-Bayes-split-kl, the results can be directly generalized to any bounded loss function by replacing PAC-Bayes-split-kl with PAC-Bayes-Empirical-Bernstein or PAC-Bayes-Unexpected-Bernstein, and deriving the corresponding analogue of Theorem 5 (which is straightforward).
  • We have shown that the bound works well with geometric split of the data, but there are many other ways to split the data which could be studied.
  • There is also a lot of space for experimentation with optimization of $\gamma_{t}$.
  • It would be interesting to study how the bound will perform in sequential learning settings, where the data arrives sequentially, and thus the partition is dictated externally.
  • There are many interesting research directions from the computational perspective. We note that for base models with linear computational complexity (e.g., neural networks) the overhead of recursion is relatively small and optimization time of Recursive PAC-Bayes is comparable to processing all data at once or in two chunks (as in data-dependent priors). For base models with superlinear computational complexity (e.g., kernel SVMs) sequential training of several small models in the recursion may actually be cheaper than training a big model based on all the data. Moreover, since the bound in Theorem 5 holds for any sequence of distributions $\pi_{0}^{*},\pi_{1}^{*},\dots,\pi_{T}^{*}$, the optimization in equation (5) is allowed to be approximate. Considering that the improvement of the bounds and the test loss relative to prior work was very significant, there is space to look at the trade-off between statistical power and computational complexity. Namely, it may potentially be possible to relax the approximation of $\arg\min$ in equation (5) to gain computational speed-up at the cost of only a small compromise on the bounds and test losses.
  • We note that it is possible to start the recursion at $\pi_{0}$. Namely, it is possible to use, for example, Theorem 2 to bound $\mathbb{E}_{\pi_{0}}[L(h)]$ using all the data, and apply the recursive decomposition (2) starting from $\pi_{1}$. Whether this would yield an advantage relative to starting the recursion at $\pi_{1}$, as we did, remains to be studied.
@@ Appendix A Illustrations
Figure 1: Evolution of PAC-Bayes. The figure shows how data are used by different PAC-Bayes approaches. Dark yellow shows data used directly for optimization of the indicated quantities. Light yellow shows data involved indirectly through dependence on the prior. Light green shows data used for estimation of the indicated quantities. In Recursive PAC-Bayes data are released and used sequentially chunk-by-chunk, as indicated by the dashed lines. For example, in the $T=4$ case $\mathbb{E}_{\pi_{1}}[L(h)]$ is first evaluated on $S_{1}$ to construct $\pi_{1}$ and $\gamma_{2}$, then in the first recursion step on $S_{1}\cup S_{2}$, in the second step on $S_{1}\cup S_{2}\cup S_{3}$, and in the last step on all $S$.
Figure 2: Recursive Decomposition into Three Terms. The figure illustrates recursive decomposition of $\mathbb{E}_{\pi_{3}}[L(h)]$ into three terms based on equation (2), and a geometric data split, as used in our experiments. The bottom line illustrates which data are used for construction of which distribution: $S_{1}$ for $\pi_{1}$; $S_{2}$ for $\pi_{2}$; and $S_{3}$ for $\pi_{3}$. The brackets above the data show which data are used for computing PAC-Bayes bounds for which term: $S_{1}\cup S_{2}\cup S_{3}$ for $\mathbb{E}_{\pi_{1}}[L(h)]$; $S_{2}\cup S_{3}$ for $\mathbb{E}_{\pi_{2}}[L(h)-\gamma_{2}\mathbb{E}_{\pi_{1}}[L(h^{\prime})]]$; and $S_{3}$ for $\mathbb{E}_{\pi_{3}}[L(h)-\gamma_{3}\mathbb{E}_{\pi_{2}}[L(h^{\prime})]]$. Note that a direct computation of a PAC-Bayes bound on $\mathbb{E}_{\pi_{3}}[L(h)]$ would have only allowed to use the data in $S_{3}$, as shown by the black dashed line. The figure illustrates that recursive decomposition provides more efficient use of the data. We also note that initially we start with poor priors, and so the $\operatorname{KL}(\pi_{t}\|\pi_{t-1})$ term for small $t$ is expected to be large, but this is compensated by a small multiplicative factor $\prod_{i=t+1}^{T}\gamma_{i}$ and availability of a lot of data $\bigcup_{i=t}^{T}S_{i}$ for computing the PAC-Bayes bound. For example, $\mathbb{E}_{\pi_{1}}[L(h)]$ is multiplied by $\gamma_{3}\gamma_{2}$ and we can use all the data for computing a PAC-Bayes bound on this term. By the time we reach higher $t$, the priors $\pi_{t-1}$ get better, and the $\operatorname{KL}(\pi_{t}\|\pi_{t-1})$ term in the bounds gets much smaller, and additionally the bounds benefit from the small variance of the excess loss. With geometric split of the data, we use little data to quickly move $\pi_{t}$ to a good region, and then we still have enough data for a good estimation of the later terms, like $\mathbb{E}_{\pi_{3}}[L(h)-\gamma_{3}\mathbb{E}_{\pi_{2}}[L(h^{\prime})]]$.
@@ B.3 Other details in the experiments
The methods in comparisons are trained and evaluated using the procedure described in Section 2 and visually illustrated in Figure 1. We will provide some further details for each method later in the following. For all methods in comparison, we apply the optimization and evaluation method described in Section 5.1. For the approximation described in Section 5.1.1, we set the parameters $c_{1}=c_{2}=5$. The lower bound for the prediction $p_{\min}=1e-5$. The $\delta$ in our bound and all the other methods is selected to be $\delta=0.025$. As mentioned in Section 5.1.2, we use the PAC-Bayes-classic bound by McAllester in replacement of PAC-Bayes-$\operatorname{kl}$ when doing optimization. Note that for all methods, we also have to estimate the empirical loss of the posterior $\mathbb{E}_{\pi}[\cdot]$ described in Section 5.1.3. We also allocate the budget for the union bound for the estimation such that these estimations in the bound are controlled with probability at least $1-\delta^{\prime}$, where we chose $\delta^{\prime}=0.01$. Therefore, the ultimate bounds for all methods hold with probability at least $1-\delta-\delta^{\prime}$. Note that we do not consider such bounds during optimization but only when estimating the bounds.
@@
In the training process of all methods in our experiments, we set the batch size to 250, the number of training epochs to 200, and use stochastic gradient descent with a learning rate of 0.005 and a momentum of 0.95.
@@ B.4.1 Recursive PAC-Bayes
We report the additional results of Recursive PAC-Bayes on MNIST with $T=2$ in Table 4, $T=4$ in Table 5, and $T=6$ in Table 6. We report Recursive PAC-Bayes on Fashion MNIST with $T=2$ in Table 7, $T=4$ in Table 8, and $T=6$ in Table 9.
@@
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .024 (3e-5) .370 (1e-3) .254 (2e-3)
2 30000 .013 (3e-3) .024 (1e-4) .136 (3e-3) .321 (3e-3) .139 (3e-3)
@@
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .023 (8e-5) .374 (1e-3) .258 (1e-3)
2 52500 -4e-4 (1e-3) .025 (3e-4) .118 (1e-3) .305 (1e-3) .126 (2e-3)
3 45000 .053 (1e-3) .002 (9e-5) .087 (2e-3) .240 (2e-3) .114 (1e-3)
4 30000 .054 (1e-3) .001 (2e-5) .083 (1e-3) .203 (8e-4) .109 (1e-3)
Table 6: Insight into the training process of the Recursive PAC-Bayes for $T=6$ on MNIST.
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .019 (7e-5) .425 (1e-3) .311 (3e-3)
2 58125 -0.013 (1e-3) .032 (6e-4) .128 (2e-3) .341 (3e-3) .139 (1e-3)
3 56250 .050 (1e-3) .003 (1e-4) .093 (6e-4) .264 (1e-3) .117 (2e-3)
4 52500 .051 (1e-3) .001 (6e-5) .080 (9e-4) .212 (5e-4) .108 (2e-3)
5 45000 .051 (1e-3) 9e-4 (3e-5) .076 (2e-3) .182 (1e-3) .104 (6e-4)
6 30000 .049 (1e-3) 7e-4 (3e-5) .074 (1e-3) .166 (1e-3) .101 (1e-3)
@@
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .011 (3e-5) .466 (1e-3) .389 (5e-3)
2 30000 .064 (3e-3) .013 (2e-4) .171 (4e-3) .404 (3e-3) .266 (5e-3)
@@
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 .011 (1e-4) .476 (4e-3) .397 (6e-3)
2 52500 .032 (6e-4) .017 (9e-4) .147 (2e-3) .386 (3e-3) .240 (4e-3)
3 45000 .100 (3e-3) 3e-3 (1e-4) .138 (5e-3) .331 (4e-3) .222 (5e-3)
4 30000 .095 (2e-3) 7e-4 (5e-5) .128 (2e-3) .293 (1e-3) .213 (3e-3)
$t$ $n^{\mathrm{val}}_{t}$ $\mathbb{E}_{\pi_{t}}[\hat{F}_{\gamma_{t}}(h,U^{\mathrm{val}}_{t}\circ\hat{\pi}_{t-1})]$ $\frac{\operatorname{KL}(\pi_{t}^{*}\|\pi_{t-1}^{*})}{n^{\mathrm{val}}_{t}}$ $\mathcal{E}_{t}(\pi_{t}^{*},\gamma_{t})$ $B_{t}(\pi^{*}_{t})$ Test 0-1
1 60000 9e-3 (8e-5) .534 (4e-3) .462 (5e-3)
2 58125 .013 (6e-3) .023 (1e-3) .151 (4e-3) .418 (5e-3) .254 (6e-3)
3 56250 .091 (3e-3) 3e-3 (4e-4) .141 (2e-3) .350 (3e-3) .223 (1e-3)
4 52500 .090 (2e-3) 9e-4 (9e-5) .121 (1e-3) .296 (1e-3) .207 (1e-3)
5 45000 .090 (1e-3) 6e-4 (3e-5) .117 (2e-3) .265 (2e-3) .199 (2e-3)
6 30000 .093 (1e-3) 5e-4 (2e-5) .122 (1e-3) .255 (1e-3) .198 (1e-3)
@@ B.4.4 Data-informed priors + excess loss
We report the additional results of data-informed priors + excess loss (Mhammedi et al., 2019) on MNIST and Fashion MNIST in Table 12 and 13. As described earlier in Section 2, Section 5, and Section B.3, we evaluate the bound using $S_{2}$ that is independent of the data-informed prior $\pi_{1}$ and the reference prediction rule $h^{*}$. The bound is composed of two parts: a bound on the excess loss of $\rho$ with respect to $h^{*}$ (Excess bound) and a single hypothesis bound on $h^{*}$ ($h^{*}$ bound). We report the two components of the bound in Table 12. We provide further details to compute these bounds from the losses of their corresponding quantities in Table 13.